# Astrophysical Implications of Eccentricity in Gravitational Waves from Neutron Star-Black Hole Binaries

### v1

@@ Abstract
The gravitational-wave signal from the neutron star-black hole (NSBH) merger GW200105 is consistent with this binary having significant orbital eccentricity close to merger. This raises the question of how an eccentric NSBH might form. Compact object binaries that form via isolated binary star evolution should radiate away any orbital eccentricity long before their gravitational-wave signal enters the sensitive frequency range of the LIGO-Virgo-KAGRA detector network. Meanwhile, dynamical environments—which can be conducive to mergers on eccentric orbits—produce very few NSBHs. Here, we focus on a formation channel that efficiently produces NSBHs with both misaligned spins and significant eccentricity close to merger: isolated field triples. We estimate the minimum measurable eccentricity of NSBHs at $10$ Hz orbit-averaged gravitational-wave frequency, $e_{\mathrm{min},10}$, finding that for GW200105, GW200115, and GW230529-like systems, $e_{\mathrm{min},10}$ is $\mathcal{O}(0.01)$. For a GW190814-like unequal-mass binary with significant higher-order mode content, $e_{\mathrm{min},10}=0.003$; this is an order of magnitude lower than when higher-order modes are not present. For dominant-mode signals from eccentric binaries with $m_{2}=1.5$ M and a range of total masses from $3\,{\rm M_{\odot}}\leq M\leq 50\,\rm M_{\odot}$, we find $0.008\leq e_{\mathrm{min},10}\leq 0.022$. The relationship between $M$ and $e_{\mathrm{min},10}$ is linear when the binaries are non-spinning. When the binaries are maximally spin-precessing, $e_{\mathrm{min},10}$ decreases as mass ratio becomes more unequal. We estimate the sensitivity of a quasi-circular templated search to a population of NSBH mergers from field triples, finding that we recover only $27\%$ of our simulated population. Finally, we show that if $\sim\nicefrac{{1}}{{3}}$ of present NSBH detections are measurably eccentric, then all of them are consistent with an isolated field triple origin.
I Introduction
@@
The LIGO-Virgo-KAGRA (LVK) Collaboration [LVK] has reported the confident detection of three neutron star-black hole (NSBH) mergers via their gravitational-wave (GW) signals. GW200105 and GW200115 [NSBHs], reported during the LVK’s third observing run (O3), have primary component masses consistent with being black holes (BHs), and secondary masses consistent with being neutron stars (NSs). GW230529 [GW230529], reported during the fourth observing run (O4), has an NS secondary and a primary with an ambiguous mass of $2.5\lesssim m_{1}\lesssim 4.5$ M, which resides within the “lower mass gap” between NS masses and BH masses; GW230529 is thought to most likely be a low-mass NSBH, given the theoretical Tolman-Oppenheimer-Volkoff (TOV) upper limit to NS masses ($M_{\mathrm{TOV}}=2.25^{+0.08}_{-0.07}\rm M_{\odot}$ [2024PhRvD.109d3052F]; see also [1939PhRv...55..364T, PhysRev.55.374, 1996ApJ...470L..61K, 2018ApJ...852L..25R]). Other candidates are consistent with being of NSBH origin, although with lower probability. GW190814 [GW190814], detected in O3, is an ambiguous merger with a secondary residing in the lower mass gap; GW190814, again due to the TOV mass limit, is thought most likely to be a binary BH merger. There are six lower-significance or ambiguous NSBH candidates in the second observing run (O2) and O3 [GWTC-2, GWTC-2-1, GWTC-3]. In the engineering run preceding O4, the NSBH candidate GW230518_125908 was detected [GWTC-4].
@@
Renewed interest in one of these NSBHs, GW200105, has recently arisen due to follow-up studies finding evidence for measurable orbital eccentricity in this GW signal [Fei2024, 2025arXiv250315393M, planas2025, Kacanja2025, 2025arXiv250812460J, 2025arXiv250926152T]. The Bayes factor in favour of the eccentric hypothesis over the quasi-circular hypothesis is low: for example, planas2025 find $\log_{10}$ Bayes factors from $\approx 0.1$ to $1.2$ depending on prior settings, indicating negligible to mild preference of the eccentric hypothesis. Nonetheless, these findings motivate further investigation into the formation, detection, and identification of eccentric NSBHs.
@@
In general, an NSBH merger may form as a result of the isolated evolution of a massive stellar binary [e.g., 1998A&A...332..173P, 2002ApJ...572..407B, 2012ApJ...759...52D, 2021ApJ...920L..13B, 2024A&A...683A.144X, 2022MNRAS.513.5780C], in which case the binary would merge with undetectable eccentricity [Peters1964, 2002ApJ...572..407B]. Alternatively, the binary may form via a pathway that allows it to retain significant eccentricity close to merger. For binary BHs, detectable eccentricities at a GW frequency of $10$ Hz, $e_{10}$, are possible when formation and merger occurs in densely-populated environments like globular or nuclear star clusters and active galactic nuclei (AGN) [e.g., 2018PhRvL.120o1101R, 2019ApJ...871...91Z, 2023:Chattopadhyay:NSCs, Tagawa2021_ecc]. However, NSBH mergers from globular cluster environments are not expected to contribute significantly to LVK observations: the predicted local merger rate densities $\lesssim 0.17$ Gpc3 yr-1[2013MNRAS.428.3618C, Bao-Minh2020, 2020ApJ... [^8] L..10Y] are low compared to the NSBH merger rate inferred from GW observations, 9.19.1-$84$ Gpc3 yr-1[GWTC-4-astro-pop]. Moreover, the few NSBH binaries that could be produced in these environments do not retain measurable eccentricities close to merger [ArcaSedda2020]. NSBHs can be formed efficiently in young star clusters, with rates $\sim 28$ Gpc3 yr-1, but are ejected and merge in the field, leading them to have negligible eccentricities at merger [2020MNRAS.497.1563R]. Detectably-eccentric NSBHs may be produced in nuclear clusters, but, similarly, local merger rate densities are predicted to be very low, $\lesssim 0.10$ Gpc3 yr-1 [2019MNRAS.488...47F]. Rates of NSBH mergers in AGN are potentially high and consistent with observed LVK rates, but their eccentricities are uncertain [2020MNRAS.498.4088M].
@@
An NSBH that is observed with detectable close-to-merger eccentricity may instead result from the evolution of isolated hierarchical triples in the galactic field, where the tertiary companion can drive large-amplitude von Zeipel-Kozai-Lidov (ZKL) oscillations of the NSBH eccentricity in the inner binary [Zeipel1910, Kozai1962, Lidov1962, Naoz2016]. As a result, efficient gravitational-wave emission at highly eccentric pericentre passages shrinks the orbit of the NSBH, effectively decouples it from further perturbation from the tertiary, and can lead to an inspiral and merger with residual eccentricity. Owing to the large abundance of massive stellar triples in the field [Moe2017, Offner2023], previous studies find that this process can efficiently produce NSBH mergers at rates consistent with or higher than that inferred from LVK observations [FragioneLoeb:2019:triples, HamersThompson:2019:triples, StegmannKlencki2025]. Thus, residual eccentricities $10^{-4}$ at $10$ Hz, and spins that are misaligned with the orbital angular momentum of the binary, are characteristic of NSBHs formed through the field evolution of triple stars [StegmannKlencki2025].
@@
Current LVK GW searches are inefficient for detecting inspiral-dominated binaries with significant eccentricities, with the loss of sensitivity worsening for higher eccentricities and lower masses [2010PhRvD..81b4007B, 2024:Divyajyoti:blind-spots, 2024PhRvD.110d4013G, Bhaumik:2025:cWBsearchsensitivity]. LVK search methods are either modelled, using quasi-circular aligned-spin waveform templates [e.g., MBTA, PyCBC, GSTLAL, SPIIR], or unmodelled, looking for correlated excess power in the data [2016PhRvD..93d2004K, 2021SoftX..1400678D].
@@
Highly-eccentric merger-dominated signals with minimal inspiral may be detected in a burst search [see, e.g., 2025arXiv250712374L, and references therein], while low-eccentricity merger-dominated signals may be picked up in a quasi-circular modelled search due to the close match with a quasi-circular aligned-spin template in the merger. Meanwhile, modelled searches are less sensitive to lower- and more unequal-mass binaries if eccentricity is excluded from search templates; such signals are also less likely to be flagged during a burst search due to the relatively lower-powered merger [e.g., 2010PhRvD..81b4007B, 2013PhRvD..87l7501H, 2024:Divyajyoti:blind-spots]. Even when both burst and modelled searches are used, one loses a significant fraction of eccentric sources [2024PhRvD.110d4013G, 2024PhRvD.110d4013G]. Thus, if there are many NSBH mergers with non-negligible orbital eccentricity—as predicted from field triple evolution—then there is a high risk of not detecting them in LVK searches. Since injection campaigns used to estimate selection effects also do not include eccentricity, the NSBH merger rate inferred from GW observations may thus be artificially skewed to lower values due to the loss of search sensitivity to eccentric NSBHs.
@@
Several modelled searches targeting inspiral-dominated eccentric systems have been developed [e.g., 2020ApJ...890....1N, 2021ApJ...915...54N, 2025PhRvD.111j3018D, 2025arXiv250805018W]. 2025PhRvD.111j3018D specifically target aligned-spin eccentric NS-containing binaries, requiring the signal to trigger a detection in more than one instrument. Their template bank covers $e_{22,10}\leq 0.46$, where the $22,10$ subscript indicates that the eccentricity is measured when the dominant $l=m=2$ ($22$) mode GW frequency reaches $10$ Hz. Since GW200105 was only detected in LIGO Livingston (although Virgo data was also analysed), it was not recovered by this search. This search is also restricted to detector-frame total masses of $\leq 10$ M, which, in combination with the eccentricity constraint, puts $\gtrsim 80\%$ of the NSBH masses and eccentricities predicted by StegmannKlencki2025 outside of the searched range even before considering the loss of search sensitivity due to the misaligned spins expected from triples. In an aligned-spin, inspiral-only targeted search for eccentric NSBHs in O3, Phukon:2025:inprep recover GW200105 at higher significance than in an equivalent quasi-circular search, with no other significant candidates. However, this search restricts to detector-frame primary masses $5\leq m_{1}\leq 15$ M and eccentricities $e_{22,20}\leq 0.2$ (corresponding to $e_{22,10}\lesssim 0.35$), thereby excluding $\gtrsim 72\%$ of the parameter space predicted by StegmannKlencki2025 before considering misaligned spins. Precessing-spin searches have also been developed specifically for NSBHs, revealing that current LVK search methods may be artificially biased against detecting this flavour of binary due to neglecting misaligned spins [2025arXiv250309773H].
@@
Long inspirals are both a blessing and a curse when it comes to detecting and measuring eccentricity. One the one hand, eccentric inspirals are less likely to be detected by quasi-circular searches, since the mismatch between an eccentric signal and a circular template leads to a reduction in SNR and astrophysical probability. On the other hand, precisely because of this larger overall difference to a quasi-circular signal, they increase our ability to measure eccentricity if analysed using an eccentric waveform model: the eccentricity of the signal becomes easier to measure if more cycles are in-band [e.g., 2020CQGra..37v5015M, 2020MNRAS.496L..64R, 2023:Romero-Shaw:Ecc-or-precc]. Furthermore, systems observed earlier in their evolution are more likely to retain measurable eccentricity, as they have undergone less GW-driven circularization. The source property that has the dominant influence on signal duration is the mass: lower-mass systems spend longer in-band because they merge at higher frequencies. Additionally, systems with more unequal masses have more cycles in-band. Accordingly, lower- and more unequal-mass binaries, such as NSBHs, can also be expected to have lower minimum detectable eccentricities.
@@
Similarly, GW signals that contain higher-order modes (HMs) may have reduced detectability due to searches using only dominant-mode templates, but also increase the information extractable from the signal. HMs are most prominent in GW signals from unequal-mass binaries, making unequal-mass NSBHs possible HM candidates. While none of the three confident NSBH mergers detected so far show signs of strong HM content, one possible NSBH candidate—GW190814—does contain HMs [GW190814, 2023PhRvD.107h4026G]. GW190814 originated from an unequal-mass compact binary merger, with source-frame masses $m_{1}\approx 23$ M and $m_{2}\approx 2.6$ M. The secondary mass of GW190814 lies in the “lower mass gap”, and its identity—NS or BH—is ambiguous. In GW signals from eccentric binaries, each mode present in the signal acquires overtones. In other words, signals with HMs contain more information about the eccentricity of the binary. As we demonstrate, this can enable the measurement of lower eccentricities than is possible when only the dominant mode is considered.
@@
In this paper, we estimate the minimum detectable eccentricity for NSBH-like mergers. For nonspinning primaries, this correlates with total mass, and tends to be $\mathcal{O}(0.01)$ for NSBH-like binaries with network matched-filter SNRs of $\rho\approx 24$. If the binary exhibits significant spin-induced precession in addition to eccentricity, the measurability of eccentricity improves for higher-mass binaries. This is because binaries with more unequal mass ratios have more spin-induced precession cycles in-band, which are altered by the presence of eccentricity. If higher-order modes are present in the signal, the minimum detectable eccentricity can be $\approx 1$ order of magnitude lower. We assess the detectability of a population of NSBH mergers produced in field triples, finding that the majority are not detected due to their relatively high eccentricities. Assuming a quasi-circular modelled search only, we miss $73\%$ of the signals from this population. By modelling the detection of NSBH mergers as a Poisson process, we find that if GW200105 was formed in an isolated field triple, then all NSBHs observed so far may have evolved via the same mechanism.
@@
We introduce our methodology in Section II, describe our results and their astrophysical implications in Section III, and conclude in Section IV.
II Methods

### v2

@@ Abstract
The gravitational-wave signal from the neutron star-black hole (NSBH) merger GW200105 is consistent with this binary having significant orbital eccentricity close to merger. This raises the question of how eccentric NSBHs form. Compact binaries that evolve in isolation radiate away any orbital eccentricity long before their gravitational-wave signal enters the sensitive frequency range of the LIGO-Virgo-KAGRA detector network. Meanwhile, dynamical environments—which can be conducive to mergers on eccentric orbits—produce very few NSBHs. Here, we focus on a formation channel that efficiently produces NSBHs with both misaligned spins and significant eccentricity close to merger: isolated field triples. We estimate the minimum measurable eccentricity of NSBHs at $10$ Hz orbit-averaged gravitational-wave frequency, $e_{\mathrm{min},10}$, finding that for GW200105, GW200115, and GW230529-like systems, $e_{\mathrm{min},10}$ is $\mathcal{O}(0.01)$. For a GW190814-like unequal-mass binary with significant higher-order mode content, $e_{\mathrm{min},10}=0.003$; this is an order of magnitude lower than when higher modes are not present. For dominant-mode signals from binaries with $m_{2}=1.5$ M and total masses from $3\,{\rm M_{\odot}}\leq M\leq 50\,\rm M_{\odot}$, we find $0.008\leq e_{\mathrm{min},10}\leq 0.022$. The relationship between $M$ and $e_{\mathrm{min},10}$ is linear when the binaries are non-spinning. When the binaries are maximally spin-precessing, $e_{\mathrm{min},10}$ decreases as mass ratio becomes more unequal. We estimate the sensitivity of a quasi-circular aligned-spin templated search to NSBH mergers from field triples, finding that we recover only $46\%$ of systems that would have been detected with a search containing the full physics of the injected population. Finally, we show that if $\sim\nicefrac{{1}}{{3}}$ of present NSBH detections are measurably eccentric, then $\geq 40\%$ are consistent with an isolated field triple origin.
keywords:
transients: black hole - neutron star mergers – gravitational waves – stars: neutron – stars: black holes
pubyear: 2026pagerange: Astrophysical Implications of Eccentricity in Gravitational Waves from Neutron Star-Black Hole BinariesB
1 Introduction
@@
The LIGO-Virgo-KAGRA (LVK) Collaboration (Abbott and others, 2018) has reported the confident detection of three neutron star-black hole (NSBH) mergers via their gravitational-wave (GW) signals. GW200105 and GW200115 (Abbott and others, 2021a), reported during the LVK’s third observing run (O3), have primary component masses consistent with being black holes (BHs), and secondary masses consistent with being neutron stars (NSs). GW230529 (Abac and others, 2024), reported during the fourth observing run (O4), has an NS secondary and a primary with an ambiguous mass of $2.5\lesssim m_{1}\lesssim 4.5$ M, which resides within the “lower mass gap” between NS masses and BH masses; GW230529 is thought to most likely be a low-mass NSBH, given the theoretical Tolman-Oppenheimer-Volkoff (TOV) upper limit to NS masses ($M_{\mathrm{TOV}}=2.25^{+0.08}_{-0.07}\rm M_{\odot}$ (Fan et al., 2024); see also (Tolman, 1939; Oppenheimer and Volkoff, 1939; Kalogera and Baym, 1996; Rezzolla et al., 2018)). Other candidates are consistent with being of NSBH origin, although with lower probability. GW190814 (Abbott and others, 2020), detected in O3, is an ambiguous merger with a secondary residing in the lower mass gap; GW190814, again due to the TOV mass limit, is thought most likely to be a binary BH merger. There are six lower-significance or ambiguous NSBH candidates in the second observing run (O2) and O3 (Abbott and others, 2021b; Abbott et al., 2024; Abbott and others, 2023). In the engineering run preceding O4, the NSBH candidate GW230518_125908 was detected (The LIGO Scientific Collaboration et al., 2025a).
@@
Renewed interest in one of these NSBHs, GW200105, has recently arisen due to follow-up studies finding evidence for measurable orbital eccentricity in this GW signal (Fei and Yang, 2024; Morras et al., 2025b; de Lluc Planas et al., 2025; Kacanja et al., 2025b; Jan et al., 2025; Tiwari et al., 2025). The Bayes factor in favour of the eccentric hypothesis over the quasi-circular hypothesis is low: for example, de Lluc Planas et al. (2025) find $\log_{10}$ Bayes factors from $\approx 0.1$ to $1.2$ depending on prior settings, indicating negligible to mild preference of the eccentric hypothesis. Nonetheless, these findings motivate further investigation into the formation, detection, and identification of eccentric NSBHs.
@@
In general, an NSBH merger may form as a result of the isolated evolution of a massive stellar binary (e.g., Portegies Zwart and Yungelson, 1998; Belczynski et al., 2002; Dominik et al., 2012; Broekgaarden and Berger, 2021; Xing et al., 2024; Chattopadhyay et al., 2022), in which case the binary would merge with undetectable eccentricity (Peters, 1964; Belczynski et al., 2002). Alternatively, the binary may form via a pathway that allows it to retain significant eccentricity close to merger. For binary BHs, detectable eccentricities at a GW frequency of $10$ Hz, $e_{10}$, are possible when formation and merger occurs in densely-populated environments like globular or nuclear star clusters and active galactic nuclei (AGN) (e.g., Rodriguez et al., 2018; Zevin et al., 2019; Chattopadhyay et al., 2023; Tagawa et al., 2021). However, NSBH mergers from globular cluster environments are not expected to contribute significantly to LVK observations: the predicted local merger rate densities $\lesssim 0.17$ Gpc3 yr-1(Clausen et al., 2013; Hoang et al., 2020; Ye et al., 2020) are low compared to the NSBH merger rate inferred from GW observations, 9.19.1-$84$ Gpc3 yr-1(The LIGO Scientific Collaboration et al., 2025b). Moreover, the few NSBH binaries that could be produced in these environments do not retain measurable eccentricities close to merger (Arca Sedda, 2020). NSBHs can be formed efficiently in young star clusters, with rates $\sim 28$ Gpc3 yr-1, but are ejected and merge in the field, leading them to have negligible eccentricities at merger (Rastello et al., 2020). Detectably-eccentric NSBHs may be produced in nuclear clusters, but, similarly, local merger rate densities are predicted to be very low, $\lesssim 0.10$ Gpc3 yr-1 (Fragione et al., 2019). Rates of NSBH mergers in AGN are potentially high and consistent with observed LVK rates, but their eccentricities are uncertain (McKernan et al., 2020).
@@
An NSBH that is observed with detectable close-to-merger eccentricity may instead result from the evolution of isolated hierarchical triples in the galactic field, where the tertiary companion can drive large-amplitude von Zeipel-Kozai-Lidov (ZKL) oscillations of the NSBH eccentricity in the inner binary (Zeipel, 1909; Kozai, 1962; Lidov, 1962; Naoz, 2016). As a result, efficient gravitational-wave emission at highly eccentric pericentre passages shrinks the orbit of the NSBH, effectively decouples it from further perturbation from the tertiary, and can lead to an inspiral and merger with residual eccentricity. Owing to the large abundance of massive stellar triples in the field (Moe and Di Stefano, 2017; Offner et al., 2023), previous studies find that this process can efficiently produce NSBH mergers at rates consistent with or higher than that inferred from LVK observations (Fragione and Loeb, 2019; Hamers and Thompson, 2019; Stegmann and Klencki, 2025). Thus, residual eccentricities $10^{-4}$ at $10$ Hz, and spins that are misaligned with the orbital angular momentum of the binary, are characteristic of NSBHs formed through the field evolution of triple stars (Stegmann and Klencki, 2025).
@@
Current LVK GW searches are inefficient for detecting inspiral-dominated binaries with significant eccentricities, with the loss of sensitivity worsening for higher eccentricities and lower masses (Brown and Zimmerman, 2010; Divyajyoti et al., 2024; Gadre et al., 2024; Bhaumik et al., 2025). LVK search methods are either modelled, using quasi-circular aligned-spin waveform templates (e.g., Alléné et al., 2025; Dal Canton et al., 2021; Joshi et al., 2025; Chu et al., 2020), or unmodelled, looking for correlated excess power in the data (Klimenko et al., 2016; Drago et al., 2021).
@@
Highly-eccentric merger-dominated signals with minimal inspiral may be detected in a burst search (see, e.g., LIGO Scientific Collaboration et al., 2025, and references therein), while low-eccentricity merger-dominated signals may be picked up in a quasi-circular modelled search due to the close match with a quasi-circular aligned-spin template in the merger. Meanwhile, modelled searches are less sensitive to lower- and more unequal-mass binaries if eccentricity is excluded from search templates; such signals are also less likely to be flagged during a burst search due to the relatively lower-powered merger (e.g., Brown and Zimmerman, 2010; Huerta and Brown, 2013; Divyajyoti et al., 2024). Even when both burst and modelled searches are used, one loses a significant fraction of eccentric sources (Gadre et al., 2024). Thus, if there are many NSBH mergers with non-negligible orbital eccentricity—as predicted from field triple evolution—then there is a high risk of not detecting them in LVK searches. Since injection campaigns used to estimate selection effects also do not include eccentricity, the NSBH merger rate inferred from GW observations may thus be artificially skewed to lower values due to the loss of search sensitivity to eccentric NSBHs.
@@
Several modelled searches targeting inspiral-dominated (i.e., long duration) eccentric signals have been developed (e.g., Nitz et al., 2020; Nitz and Wang, 2021; Dhurkunde and Nitz, 2025; Wang and Nitz, 2025). Dhurkunde and Nitz (2025) specifically target aligned-spin eccentric NS-containing binaries, requiring the signal to trigger a detection in more than one instrument. Their template bank covers $e_{22,10}\leq 0.46$, where the $22,10$ subscript indicates that the eccentricity is measured when the dominant $l=m=2$ ($22$) mode GW frequency reaches $10$ Hz. Since GW200105 was only detected in LIGO Livingston (although Virgo data was also analysed), it was not recovered by this search. This search is also restricted to detector-frame total masses of $\leq 10$ M, which, in combination with the eccentricity constraint, puts $\gtrsim 80\%$ of the NSBH masses and eccentricities predicted by Stegmann and Klencki (2025) outside of the searched range even before considering the loss of search sensitivity due to the misaligned spins expected from triples. In an aligned-spin, inspiral-only targeted search for eccentric NSBHs in O3, Phukon et al. (2025) recover GW200105 at higher significance than in an equivalent quasi-circular search, with no other significant candidates. However, this search restricts to detector-frame primary masses $5\leq m_{1}\leq 15$ M and eccentricities $e_{22,20}\leq 0.2$ (corresponding to $e_{22,10}\lesssim 0.35$), thereby excluding $\gtrsim 72\%$ of the parameter space predicted by Stegmann and Klencki (2025) before considering misaligned spins. Precessing-spin searches have also been developed specifically for NSBHs, revealing that current LVK search methods may be artificially biased against detecting this flavour of binary due to neglecting misaligned spins (Harry and Hoy, 2025).
@@
Long inspirals are both a blessing and a curse when it comes to detecting and measuring eccentricity. One the one hand, eccentric inspirals are less likely to be detected by quasi-circular searches, since the mismatch between an eccentric signal and a circular template leads to a reduction in SNR and astrophysical probability. On the other hand, precisely because of this larger overall difference to a quasi-circular signal, they increase our ability to measure eccentricity if analysed using an eccentric waveform model: the eccentricity of the signal becomes easier to measure if more cycles are in-band (e.g., Moore and Yunes, 2020; Romero-Shaw et al., 2020b, 2023). Furthermore, systems observed earlier in their evolution are more likely to retain measurable eccentricity, as they have undergone less GW-driven circularization. The source property that has the dominant influence on signal duration is the mass: lower-mass systems spend longer in-band because they merge at higher frequencies. Additionally, systems with more unequal masses have more cycles in-band. Accordingly, lower- and more unequal-mass binaries, such as NSBHs, can also be expected to have lower minimum detectable eccentricities.
@@
Similarly, GW signals that contain higher-order modes (HMs) may have reduced detectability due to searches using only dominant-mode templates, but also increase the information extractable from the signal. HMs are most prominent in GW signals from unequal-mass binaries, making unequal-mass NSBHs possible HM candidates. While none of the three confident NSBH mergers detected so far show signs of strong HM content, one possible NSBH candidate—GW190814—does contain HMs (Abbott and others, 2020; Gonzalez et al., 2023). GW190814 originated from an unequal-mass compact binary merger, with source-frame masses $m_{1}\approx 23$ M and $m_{2}\approx 2.6$ M. The secondary mass of GW190814 lies in the “lower mass gap”, and its identity—NS or BH—is ambiguous. In GW signals from eccentric binaries, each mode present in the signal acquires overtones. In other words, signals with HMs contain more information about the eccentricity of the binary. As we demonstrate, this can enable the measurement of lower eccentricities than is possible when only the dominant mode is considered.
@@
In this paper, we estimate the minimum detectable eccentricity for NSBH-like mergers. For nonspinning primaries, this correlates with total mass, and tends to be $\mathcal{O}(0.01)$ for NSBH-like binaries with network matched-filter SNRs of $\rho\approx 24$. If the binary exhibits significant spin-induced precession in addition to eccentricity, the measurability of eccentricity improves for higher-mass binaries. This is because binaries with more unequal mass ratios have more spin-induced precession cycles in-band, which are altered by the presence of eccentricity. If higher-order modes are present in the signal, the minimum detectable eccentricity can be $\approx 1$ order of magnitude lower. We assess the detectability of a population of NSBH mergers produced in field triples, finding that the majority are not detected due to their relatively high eccentricities. Assuming a quasi-circular modelled search only, we detect only $46\%$ of the signals that could have been recovered using a search that contains the full physics of the population. By modelling the detection of NSBH mergers as a Poisson process, we find that if GW200105 was formed in an isolated field triple, then all NSBHs observed so far may have evolved via the same mechanism.
@@
We introduce our methodology in Section 2, describe our results and their astrophysical implications in Section 3, and conclude in Section 4.
2 Methods